### v2

# Engineering Scientific Assistants using Interactive Structured Induction of Programs

Abstract
We are interested in the construction of software that can act as scientific assistants to domain specialists. It is expected that such assistants will be needed to accelerate the identification of ways to address complex problems requiring urgent solutions. In this paper, our focus is not on a specific scientific problem, but on the software-engineering of such ‘science accelerators’. Recent developments in ‘No Code’ techniques would seem to suggest that scientist can simply hypothesise solutions simply by conversing with a large language model (LLM). However, for complex scientific problems, this seems unlikely given the current state of LLM technology. What does appear feasible is that a software engineer can use LLMs to rapidly construct programs for use by a domain-specialist, including the specialist’s requirements expressed in natural language. We propose the design of an interactive form of ‘structured’ inductive programming in which a software-engineer and an LLM collaboratively construct an ‘assistant’ for a scientific data analysis. The paper describes a simple implementation called that adapts a ‘2-way Intelligibility’ protocol to implement the interaction between the software engineer and the LLM. We test the tool on two different non-trivial scientific data analysis tasks. Specifically, we compare the system constructed by against systems constructed manually and by Low Code/No Code methods along dimensions of: (a) program performance; (b) program quality; and (c) programming effort. The results show allows a software engineer to develop better programs faster suggesting interactive structured induction can play a useful role in the rapid construction of scientific assistants.
keywords:
Structured Induction, Inductive Programming, Human-AI Collaboration1 Introduction
In The Mythical Man-Month, Fred Brooks famously identified software as perhaps the most complex of human creations [1], although the idea of programs as exceedingly complex artifacts had already been raised by Herb Simon. But well before both Brooks and Simon—focussing on the artisan rather than the artifact—Dijkstra stated: “As the power of available computing increases, the problems tackled by programmers have grown ever more complex, and programming itself has become a most challenging intellectual endeavour.” [2]. Today, practising software engineers may be even more sympathetic to Dijkstra’s dilemma than their counterparts of five decades ago. To date, algorithm design and program construction have remained largely resistant to significant automation. However, the advent of foundation models trained on billions of lines of code may be poised to change this situation. Large language models (LLMs) are now being used to generate code for a wide range of tasks, from routine programming (for example, form-filling and basic web design) to more open-ended activities, such as automating aspects of scientific research [3, 4, 5, 6]. Unsurprisingly, as LLMs become increasingly capable at code generation, Dijkstra resurfaces: software engineers—now armed with LLMs—are expected to construct software for problems where human prior knowledge and intuition may be limited or even absent. We are at an early stage of understanding how best to use tools like LLMs for program construction.
One approach to addressing the problem of overly-complex programs constructed by machine was proposed by Shapiro [7], in the form of structured induction [8]. Clearly motivated by the idea of structured programming the intent is that human software specialist will break down the overall task into sub-tasks. An ML technique is then used to construct automated procedures for executing the sub-task. In Shapiro’s case, the overall task is restricted to one of constructing a program for classification of data instances. Sub-tasks are identification of domain-specific concepts that are relevant to the overall classification problem (the concepts are manually identified: subsequent developments examined how these could be automatically determined [9]). Programs for performing these sub-tasks are constructed by an ML engine (a decision-tree learner). Human-labelled instances act as (partial-) specifications for the sub-tasks and the decision-tree constructed for each concept is required to be consistent with the corresponding partial specification. Shapiro shows that this structured construction of sub-programs makes the result both understandable and verifiable by a domain-specialist. Shapiro’s work can be considered a form of interactive Inductive Programming (IP: [10], although IP is not always motivated by human-understandability (or, what Djikstra calls “convincing programs”).
While Shapiro’s use of ML was restricted to constructing programs, it was restricted to conditional statements characteristic of decision-tree builders, we are now in a position to consider significantly more expressive programs through that of modern language models (LLMs). This has become possible due to their prior knowledge of programs constructs obtained from very large code-repositories like Github; and their ability to use this information to generate new programs (in fact, LLMs may prove better at generating convincing programs than at generating convincing natural language).
In this paper, we are concerned with the engineering of systems for accelerating scientific data analysis. By ‘system’ we mean not just a model for data as was done by Shapiro, but a connected set of modules that constitute an end-to-end workflow. Each module is a conceptually meaningful sub-task, implemented by a program constructed interactively using an LLM. Ideally, we would like the interaction to be directly between a domain-specialist and the LLM. However, neither the current state of LLM technology or specialists’ knowledge of their use appear to allow this and the involvement of a software engineer seems inevitable. Since the software engineer’s knowledge of the data analysis task is probably no better than the specialist’s knowledge of LLM usage, we will assume some general functional abstractions are possible, irrespective of the scientific task being addressed.111This is reasonably well accepted in business seeking to automate diverse data analysis workflows across operations. For example, the CRISP-DM process model (‘Cross-Industry Standard Process for Data Mining’ [11]) uses a workflow with the following (sub-)processes: (1) Business understanding, that identifies the analysis to be done; (2) Data understanding, that using the output identifies the data to be collected; (3) Data preparation, that constructs the data sources; (4) Modelling, that constructs models for the data; (5) Evaluation, that evaluates the models; and (6) Deployment, that incorporates the model into business practice. Each sub-process can be seen as implementing a function, using the output of the previous process as input. This allows us to use an approach to structured system design through the use of Data Flow Diagrams (DFDs) [12, 13]. In particular, each process element in a DFD represents a ‘guarded function’, with pre- and post-conditions. Unlike the formal system design, both pre- and post-conditions can include statements in natural language, and as in Inductive Programming, post-conditions can include a partial specification of input-output pairs that capture programmatic and domain-specific requirements. Sub-programs implementing the functions are constructed by traversing the graph representing the DFD. The programs are constructed by the software engineer interacting with an LLM adapting a recently proposed ‘2-way Intelligibility’ protocol to allow the engineer to ratify that the programs constructed satisfy the post-conditions. The result is a set of user-ratified programs that are combined using the connectivity information entailed by the DFD. Since the procedure combines aspects of both structured programming and Inductive Programming, we call this approach ‘Structured Inductive Programming’ (SIP).
2 Structured Inductive Programming using LLMs
We propose Structured Inductive Programming (SIP) as a form of Inductive Programming (IP). We assume the following informal specification of IP:
- Given.
(a) Prior or background knowledge, describing domain-specific and domain-independent information possibly relevant to program identification process; (b) A specification of the program to be identified; and (optionally) (c) functions and relations for enumerating elements from the (possibly infinite) set of programs
- Find.
A program that is consistent with the specifications.
Reviews of various kinds of IP can be found in [14]. In Structured Inductive Programming, we require: (i) the background knowledge (a) to contain information about the decomposition of the program into a set of sub-programs; (ii) the specification (b) to contain specifications and description for each of the sub-programs; and (iii) enumeration of programs in (c) is controlled by the execution of an interaction model in which the software engineer is required to ratify sub-programs constructed automatically. In the specific implementation we describe below, the decomposition information is represented by a form of Data Flow Diagram [13]; and automatic program construction is done by an LLM.
Example 1 (Structured Specification Using Data Flow Diagrams).
The principal requirement of background knowledge in this paper is the provision of a data flow diagram describing information flow in the program we want to construct.
Examples of DFDs of the kind we will use can be found in Figures 2, 3. Each process element contains a description (not shown in the figure for reasons of space), along with pre- and post-conditions, all specified as statements in natural language. Other elements in the diagram are data sources, data storage and edges (arrows) with labels denoting the information transferred.
Example 2 (Program Induction with an LLM).
An LLM can be used as a generator of programs from specifications. A prompt is constructed for an LLM using following function specification to create embedding models from Sequence data:
Description: Use word2vec to create embeddings of protein sequences for various combinations of k-mer: 3 and 5; context window: 5, 10 and 25; vector size: 100, 200 and 300; and training model: skip-gram.
Pre-condition: The swissprot and reduced sequences files are present in the ’reduced_sequences’ directory
Post-condition: Embedding models created and saved for later retrieval
A fragment of the code produced by the LLM is shown below:
import os
import gensim
from Bio import SeqIO
def generate_kmers(sequence, k):
Function to generate kmers
return [sequence[i:i+k] for i in
range(len(sequence) - k + 1)]
# Parameters
kmers_values = [3, 5]
window_values = [5, 10, 25]
vector_size_values = [100, 200, 300]
sg_values = [1] # Skip-gram
# Get list of files in directory ’reduced_sequences’
<...rest of the code...>
In the approach we propose, an LLM is used to construct a program for each process element specified in the DFD (examples of the actual prompts used for one of the problems we examine is in the Appendix). Each program is the result of an interactive session with a software engineer-in-the-loop, who ratifies, refutes or rejects programs using an interaction model described below.
2.1 Interactive Structured Inductive Programming
We first clarify the following assumptions underlying the approach we adopt.
First, our principal requirement is the background knowledge which contains a structured decomposition of the problem in the form of a data flow diagram (DFD). For us, DFDs are directed graphs with vertex- and edge-labels. Vertices denote one of: processes, data-sources or data-stores. A vertex-label denote the information related to the vertex. A directed edge from vertex to node denotes that that the output of is an input to . The edge-label denote the information being transferred (this can refer to some or all of the data, content, and type of information); Secondly, process vertices represent guarded functions. The vertex-label is the triple consisting of: a function-description, its pre-conditions, and post-conditions, - all in natural language. For vertex , we will denote these by and respectively; Finally, we will assume that the systems we are designing do not have any feedback loops. This translates to the DFDs being directed acyclic graphs. Thus, we will not be examining systems where a process requires data from a later process . In functional terms, the overall function being computed by the system can then be written as a sequential composition of the sub-processes.
Procedure 1 is a simple graph-traverser that examines total-orderings of processes in the DFD to find acceptable programs for each process. Programs are constructed by a human agent (here, a software engineer) interacting with an LLM to identify programs acceptable to the human. The interaction model used is an implementation of a ‘2-way Intelligibility’ protocol [15]. We describe this before presenting its implementation in Procedure .222For clarity, the implementation of shown does not contain measures taken for efficiency. Thus, as shown, may repeatedly construct a program for the same function as we explore different orderings. This is clearly not required: once we have obtained an acceptable program for a function, it is stored for re-use.
Input: : an LLM-based agent;
: a human-agent;
: background knowledge that includes
: a labelled DAG representation of the data flow diagram for
the task,
: description of the scientific task;
: an upper-bound on the number of retries to obtain a non-empty program;
: upper-bound on number of messages exchanged between and ;
: message-number after which an agent can send tags in messages
Output: , where is a set of programs implementing
the functions in , and is a program implementing the main task
The Interaction Model
We use as starting point the protocol described in [15]. It is motivated by the question of when an interaction between agents could be said to be intelligible to either. For us, the main features we use from the protocol are these:
Messages, other than the first and last, are accompanied by 1 of 4 “tags”, namely: or (the first message has an tag, and the last has ). We will continue to use these message-tags;
Although the protocol is intended for agents exchanging predictions and explanations about data instances, intelligibility is simply a property deducible from the message-tags exchanged. We will continue to use this definition of intelligibility (reproduced below).
There are two important differences to the setting in [15]. First, communication between agents is about programs and explanations about them. Secondly, the interaction is not symmetric: here the interaction is human-centred and refers to programs constructed by the machine. To accommodate for these differences, we propose the following modifications:
Messages exchanged of the form , where consists of the triple , denoting a function description, pre- and post-conditions (all in natural language). is a (possibly correct) implementation of , and is a description of the implementation steps. Other than the initial message, the component of the message from the human to the machine will be the program received from the machine;
The human initiates the interaction with the message . Thereafter human and machine exchange messages with tags determined by the tabulations in Figure 1 (a prediction or explanation of ’?’ is taken not to Match or Agree by either agent). The asymmetric role of the agents is reflected in the tabulations that show that the human agent never sends a and the machine agent never sends a ; For practical reasons, we will assume that the human-agent will only send a after receiving at least some pre-specified number of incorrect programs and explanations.
| Explanation Recd. | |||
| Program Recd. | Agree | Agree | |
| -4\@nil | Match | ||
| -4\@nil | Match | or | |
| -4\@nil | |||
| Explanation Recd. | ||||
| Program Recd. | Agree | Agree | ||
| -4\@nil | Match | or | ||
| -4\@nil | Match | or | ||
| -4\@nil | ||||
We retain the definitions of 1- and 2-way intelligibility proposed in [15]. Given the tabulation in Figure 1, it is evident that messages from human to machine will never contain ; and messages from machine to human will never contain . In practice, we can expect the LLM to revise its programs often in response to the feedback from the software engineer, we can expect that most sessions would be One-Way Intelligible for the LLM. Since no sequence can contain both a and tags, if we ensure the interaction terminates after a from the human, then the session will be One-Way Intelligible for the human according to the definitions in [15]. The implementation in Procedure enforces this additional constraint.
Input: : an LLM-based agent;
: a human agent;
: a task-specification consisting of the function-description, pre-condition, and post-condition;
: an initial context for the LLM;
: an upper-bound on the number of retries to obtain a non-empty program;
: an upper-bound on the number of messages exchanged between and ;
: message-number after which an agent can send tags in messages
Output: where is a program and is a context.
2.2 Note on Human-Generated Refutations
The procedure is at the heart of interactive structured induction. For each function to be implemented, it initiates interaction with the LLM using the sub-task description, pre- and post-conditions and task description as context. After that it repeatedly: (a) Obtains a program from the LLM; (b) Obtains a “message” from the human, that consists of a tag, a program and an explanation; (c) If the tag is then interaction halts and the program returned is . If the tag is then interaction halts and the program constructed by the LLM is returned. Otherwise, if the tag is interaction returns to (a) with a refutation of the program used to update the LLM’s context.
For reasons of space, we do not describe the procedures and . It is evident that in any multi-step interaction with the machine, the human is refuting the machine-generated program or its explanation or both. Jumping ahead, the refutation (Step 22) is obtained from the human. We show by way of examples the main kinds of refutations that occur in the experiments that follow.
Example 3 (Erroneous Program).
A prompt is constructed for an LLM using following function specification to train a deep learning model to predict 3 properties of star-formation in an astrophysics problem:
Description: Train a deeplearning model that will predict the three output parameters with lowest error;
Pre-condition: X_train, y_train, X_test, y_test datasets are available;
Post-condition: Model trained and saved. Predicted and actual values saved.
The LLM produces a program (and an explanation ) which results in a runtime error. The software engineer decides the program is incorrect, but the explanation is correct, and constructs the following message to the LLM in: Step 22: ,“ValueError: The filepath provided must end in ‘.keras‘ (Keras model format). Received: filepath=best_model.h5”
Example 4 (Incorrect Program).
A prompt is constructed for an LLM using following function specification to create reduced sequences based on various alphabet reduction techniques in a biochemistry problem:
Description: Create reduced sequences for train.csv and test.csv which contain 2 columns …Save the resultant sequences in files with appropriate naming. Print a list of the file names at the end; Pre-condition: train.csv and test.csv files are available; Post-condition: Files reduced_train.csv and reduced_test.csv for each type of alphabet reduction technique with appropriate reduction technique prefixes.
The LLM produces as output the program and an explanation which creates incorrect groupings. The software engineer decides the program is incorrect, but the explanation is correct, and constructs the following messaage in step 22: ,“Each amino acid in the group LIVFMWAY maps to say A. i.e. L maps to A, I maps to A, V maps to A and so on. Please make the changes in the code.”
3 Empirical Evaluation
3.1 Aim(s)
We intend to investigate the use of as a tool for human-guided construction of systems for scientific data-analysis. Specifically using two non-trivial scientific data-analysis tasks we test if: (a) can construct a program that correctly achieves the post-conditions specified for each task; and (b) The end-to-end solution constructed by performs at least as well as an end-to-end solution constructed: (i) manually for the same task by an experienced software engineer; and (ii) “Low code/No code” (LCNC) solutions constructed directly using an LLM or a standard program co-pilot working with an experienced software engineer.
For the comparative assessment in (b), we evaluate how the system compares along dimensions of: (i) performance; (ii) software quality; and (ii) time for construction.
3.2 Materials
3.2.1 Problems
Astrophysics (PHY). The scientific task involves numeric prediction of three star formation properties - star formation rate, dust luminosity and stellar mass, using multi-spectral data. We use data from 21 bands consisting of far and near ultra-violet, Sloan Digital Sky Survey bands, near infrared bands, WISE mid-infrared bands and the five Herschel bands along with spectroscopic redshift. The data is obtained from the GAMA catalogue [16] and post filtering gives a dataset of 76,455 galaxies.
Biochemistry (BIO). The main scientific task is to classify a sequence as antimicrobial peptide (AMP) or not. We use machine learning models using alphabet reduction and distributed vector representation for creating embeddings and classifying a sequence as AMP and non-AMP. The alphabet reduction is based on various physico-chemical properties of peptide sequences. Uniprot[17] data is used to create the embedding models for combinations of alphabet reduction techniques, k-mer, context window and vector size. The dataset for classification is created using publicly available databases with a size of 14,821 AMP and non-AMP sequences.
3.3 Background Knowledge
The primary requirement for using is to devise a two component background knowledge: (i) in the form of a decomposition of the overall task into sub-tasks and to construct a data flow diagram (DFD), in which process elements represent guarded functions. (ii) in the form of the scientific task description
DFD for PHY. The following sub-tasks (processes) were identified for the PHY problem: (P1) Load data and perform exploratory data analysis; (P2) Feature engineering; (P3) Train a Machine Learning model to predict the three output parameters; (P4) Create plots to compare the actual and predicted values. The corresponding data flow diagram is shown in Figure 2. The task description is shown in Appendix A.
DFD for BIO. The following sub-tasks were identified for the BIO task: (P1) Download data from Swissprot; (P2) Apply alphabet reduction technique on the swissprot dataset; (P3) Create embedding models from the swissprot data for the raw and alphabet reduced sequences; (P4) Load and join the positive and negative AMP datasets. This forms the classification dataset; (P5) Apply alphabet reduction technique on the AMP classification dataset; (P6) Create embeddings for the train and test sequences using the models created in P3; (P7) Classify using ML algorithms; (P8) Evaluate the models created and elect the best one. The corresponding data flow diagram is shown in Figure 3. (For reasons of space we do not show the task description for BIO)
3.4 Method
In order to evaluate the performance of interactive structured induction, we follow a straightforward method:
For PHY and BIO:
Call with the appropriate background knowledge (the DFD for the task, containing the information on the pre- and post-conditions for each process and the overall task description);
Obtain an overall program using the combination of and . This will be denoted as the program obtained with ;
Compare the program obtained with against programs constructed entirely manually by the software engineer; and against “Low Code/No Code” (LCNC) options along dimensions of: program performance, program quality, and programming effort.
We note first an important caveat. In order to reduce variability arising from different human subjects, we have selected the tasks for which the same software engineer is involved in both developing the manual solution and in the use of . However, even in this case, the comparisons have to be interpreted with caution, since there will be some ‘memory-effect’ arising from the human-subject having seen the problems earlier. Correctly, we would need to conduct matched experiments with multiple software engineers, randomly assigning the order of using and manual construction. As it stands, the results here should be seen as a pilot for examining whether such a larger-scale study is warranted.
The following additional details are relevant: (a) For all experiments, we use the following parameter setting (for and ): , and . In addition, for all calls to the LLM temperature was set to . (b) We assess performance on ‘test’ data. The details are as follows: (i) For PHY, where is the magphys[20] SED fitting value.; (ii) For BIO, the models are evaluated on accuracy, F1-score and Matthew Correlation Coefficient (MCC). For brevity, we report only on the in this paper. (c) We assess software quality along the following dimensions: (i) Logical checks using pylint [21]; (ii) Type violations obtained using mypy [22]; (iii) Cyclomatic complexity using radon [23]; (iv) Code complexity, measured by number of lines of code generated. (d) System development effort is assessed along the following dimensions: (i) Number of days taken to obtain the final version of an end-to-end solution; (ii) Number of software engineers involved in developing the solution; (iii) Person-months (PMs, obtained from (i) and (ii) above); (iv) Number of interactions: this denotes the total number of interactions (across all sub-tasks) between human and machine. This is of relevance for and the “Low Code” options (see next). (e) “No code” options we consider are: (i) LLM-0, which denotes directly asking the LLM to generate code for the problem; and (ii) CP-0, which denotes asking a program co-pilot to generate code for the problem. Both involve no additional involvement on behalf of the software engineer. “Low code” options we consider are: (i) LLM-k, which denotes an LLM constructing a program for the problem; and (ii) CP-K, which denotes a program co-pilot constructing a program for the problem. Both are given free-form feedback from the software engineer (that is, the interaction model is not used). To be comparable, the number of interactions are restricted to the number of messages sent with interactive structured induction. (f) We use Azure OpenAI GPT-4 [24] as the Large Language model in all experiments. For the experiments the machines used are as follows: (i) A 2.4 GHz Intel Core i9 machine with 64 GB 2667 MHz DDR4 RAM; (ii) A machine with Ubuntu (24.04 LTS) operating system, and hardware configuration such as: 64GB of main memory, AMD Ryzen 9 5900X 12-core processors. (g) The implementation shown accumulate the context across retries for any single program (Procedure 2) and across programs (Procedure 1). However, this usually results in exceeding the token-length restriction for the LLM. In experiments we employ a modification to Procedure 2 that obtains a summary of the context returned by the procedure. This summary is constructed by the LLM.
3.5 Results
The principal results are shown in Figure 4(a)–(c). The main findings against the manually-constructed solution can be summarised as follows:
Performance. For the regression problems in PHY, the performance of ’s solution constructed interactively by a software engineer appear to have substantially lower error than the manual solutions constructed by the same software engineer. For BIO, the errors from using are comparable to the manual effort.
Quality. For both problems, the quality of code obtained by the use of is better than the quality of code constructed manually. Specifically, the programs with are substantially more compact. The code generated by has lower complexity compared to the manual code and low code. This is true for the BIO problem too, even though the complexity value is slightly higher. This is primarily due to the increased number of combinations (18 times) of models created compared to the manual solution. It is interesting that is able to produce more complex code which is more compact and has a higher logic score.
Effort. The time taken to construct an end-to-end solution by may be faster by an order-of-magnitude than the time taken to construct the manual solution.
It is unsurprising that the “No Code” option did not produce legal programs (here “Error” means the program results in runtime errors). Results are slightly better with the “Low Code” approach. Performance is poorer, but it is possible that this may improve with longer interaction (the apparently low value of complexity, measured by lines of code, is misleading: this is likely to increase as higher performance values are required). The very low programming effort by the Low Code/No Code options is unsurprising. What is surprising though is that the programming effort for is almost the same, with significantly better performance. This suggests the effort of producing the DFD for each problem is not an unduly onerous task for a software-engineer. Finally, the results with LLM-k provide evidence for the combined effect of the use of structuring and of the incremental model. The LLM used in both and is the same, as are the number of interactions. What is different is that LLM-k does not employ the explicit structuring specified by the DFD, and the feedback does not follow the specification by the interaction model.
| Problem | Prediction | Error | |||||
| -8\@nil | Engineered | Low Code | No Code | ||||
| -8\@nil | Manual | CP-k | LLM-k | CP-0 | LLM-0 | ||
| Star formation | 0.030 | 0.164 | 0.332 | 0.357 | Error | Error | |
| PHY | Dust luminosity | 0.022 | 0.114 | 0.133 | 0.195 | Error | Error |
| Stellar mass | 0.020 | 0.058 | 0.140 | 0.149 | Error | Error | |
| BIO | AMP Class. | 0.024 | 0.021 | Error | Error | Error | Error |
| Problem | Measure | Manual | CP-k | LLM-k | CP-0 | LLM-0 | |
| Logic score (/10) | 6.69 | 4.57 | 1.88 | 1.71 | – | – | |
| PHY | Type errors | 1 | 1 | 0 | 0 | – | – |
| Complexity | 1.11 | 1.58 | 1.33 | 1.0 | – | – | |
| Lines of code | 259 | 564 | 48 | 50 | – | – | |
| Logic score (/10) | 6.32 | 3.54 | – | – | – | – | |
| BIO | Type errors | 0 | 48 | – | – | – | – |
| Complexity | 2.17 | 1.7 | – | – | – | – | |
| Lines of code | 754 | 3902 | – | – | – | – |
| Problem | Measure | Manual | CP-k | LLM-k | CP-0 | LLM-0 | |
| Days (approx) | 4 | 30 | 4 | 4 | 1 | 1 | |
| PHY | People | 1 | 2 | 1 | 1 | 1 | 1 |
| PMs (appox) | 0.13 | 2 | 0.13 | 0.13 | 0.03 | 0.03 | |
| Interactions | 13 | – | 13 | 13 | 1 | 1 | |
| Days (approx) | 10 | 60 | 5 | 4 | 1 | 1 | |
| BIO | People | 1 | 3 | 1 | 1 | 1 | 1 |
| PMs (approx) | 0.3 | 6 | 0.2 | 0.13 | 0.03 | 0.03 | |
| Interactions | 22 | – | 22 | 22 | 1 | 1 |
Taken together, the results tabulated here suggest that rapid, high quality scientific data analysis programs can be constructed by a software engineer interacting with a commodity LLM using interactive structured induction. Further, there is evidence that constructing programs for data analysis in this manner may be better than purely manual efforts, or Low Code/No Code options.
4 Related Work
Clearly drawing on ideas from structured programming, Shapiro’s work on structured induction [7] is the main inspiration for the work for this paper. To the best of our knowledge, it represents the first attempt to decompose machine-construction of a program into learning sub-programs. Shapiro’s approach was restricted to the construction of hierarchies of tree-based models, and was not controlled by any interaction model. The need for structuring ML-generated programs to assist human-understandability was also a feature of early ILP programs like Duce [25], has continued up to modern ILP systems like Deeplog [26], that attempt to structure programs automatically through invention of new predicates. ILP is one form of inductive programming , in which the representation of background knowledge, data and programs are all logic programs, and the search procedure is usually some form of discrete-optimisation. Other representations and approaches are clearly possible; see [27] and [28]. Examples of implementations using representations other than logic programs and other forms of search are Flashfill [29, 30] and DreamCoder [31]. The latter is interesting since it involves re-use of sub-programs across multiple tasks. Both cases are somewhat restricted in their use though and do not consider end-to-end data analysis tasks as we do here. DreamCoder also constructs functional programs: like with most ILP systems, this is now increasingly unfamiliar to even software engineers. Nevertheless, the broader principle of specifying the requirements of the program needed, provision of background knowledge, and conducting a search over programs are all features of inductive programming which we inherit in this paper.
Inductive programming itself is one form of program synthesis. A survey, although a little dated, but still very useful is [32]. It describes approaches to program synthesis, its applications, and subfields. The significant change since this review is the advent of LLMs. While their facility to generate natural language is now almost taken as commonplace, it is their–perhaps even better–ability to generate code that we rely on in this paper. Structured prompting to enhance the LLMs performance is shown in [33] which scales in-context learning and in [34] which uses cognitive prompting to enhance structured reasoning in language models. Relevant approaches to this topic include structured prompting methods, such as Chain of Thought [35]. In fact, the initial session specification given to the LLM in the form of function-description, pre-condition and post-condition can be viewed as a form of structured prompting. A systematic survey of prompt engineering in Large Language models can be found in [36].
Austin and colleagues [37] evaluate LLMs for synthesising short python programs from natural language description using few-shot learning with a well-designed prompt and fine-tuning. This is akin to the Low Code approach we evaluated with the free form interaction. This demonstrates the utility of human-interactive feedback. That the current state-of-the-art LLMs alone are not sufficient to construct complex programs has also been demonstrated in [38], where it is shown that integrating LLMs and enumerative synthesis performs better than just using a No Code approach.
We note that tools like Copilot [39] and AutoML [40] are clearly of relevance to the topic of this paper. As we saw in the empirical results, Copilot with human-feedback does produce programs, albeit with inferior performance. Copilot generates real time code suggestions while writing code within a development environment i.e. an interaction with the code is required. This contrasts with the use of interaction in our approach where interactions are with the LLM using natural language. Tools like AutoML, automates the process of developing machine learning models. In fact, there is nothing preventing the specification we provide to model-construction sub-task to use an engine like AutoML for constructing models (we have conducted such an experiment, and can confirm that it is possible).
We turn finally to the interaction model used here. It is our contention that the model plays an important role in regulating the feedback needed to construct the sub-programs. The idea that interaction between human and machine should follow a communication protocol, and that execution of the protocol will allow us to define properties like intelligibility of information exchanged is proposed in [15]. However, little is done in that paper to show any practical use-cases: the studies here can be seen as rectifying that shortcoming. The protocol-based approach is one way of specifying interaction: approaches based on dialogue-models are also possible (for example, see [41]). These are more open-ended and therefore harder to specify formally. The techniques proposed in this paper are relevant not just in the areas of PHY and BIO that we have tested, but across a range of topics, for example the ones specified in this report [42].
5 Conclusions
In this paper, we have focussed on engineering AI-based software tools for scientific data analysis. We believe the need for such tools will only increase as the nature of problems to be addressed becomes increasingly complex, the data become more diverse, and the need for solutions become more urgent. Our results here provide evidence that it is possible for a software-engineer to construct an end-to-end data analysis tool by working interactively with a large language model (LLM). The approach we have proposed combines time-tested principles from structured design of software systems and inductive programming with the ability of modern LLMs to generate programs given requirements in a natural language. The interaction follows an ‘intelligibility protocol’, adapted to the automated construction of programs. We view our approach as a generalised form of the original proposal by Shapiro of human-driven structured induction of programs by using machine learning. The results here provide evidence supporting the claims that interactive structured induction of programs with LLMs can be used for the rapid construction of programs that are both intelligible to the human engineer and perform capably as data-analysis tools.
The approach here can be improved in several ways. First, inductive programming systems often employ significantly more powerful search procedures than the simple breadth-first search we have used. The incorporation of utility measures over programs and their use in a heuristic search remains to be investigated. The use of data flow diagrams without feedback loops means data produced later cannot be used by functions earlier in the sequence. While this has not been necessary for the problems considered here, it precludes constructing some kinds of programs (those requiring mutual recursion, for example). The need to provide a structured breakdown of the problem does require some effort on the part of the human engineer. It is feasible that LLMs can be used to suggest a structuring. We would then see the interaction model extend to arrive additionally to an intelligible form of structuring. As it stands, we can see the structuring of the domain results in a form of structuring of the prompts according to sub-tasks. In the longer term, a facility for automated structuring may enable the construction of data analysis programs directly by a domain-specialist.
Appendix A Prompts used for PHY
Here we show the initial prompts given to the LLM for each of the sub-tasks in the PHY problem. This is provided at the start of each session for each sub-task. The function-description given to the procedure is a combination of the task description and the sub-task specification combined in this manner:
" + While the whole task description is given above, in this current session let us focus only on the users given sub-task: ".
We describe below the task description given for the PHY problem, followed by each session specification which consists of the sub-task specification, pre-condition and the post-condition. This prompt is used at the start of the session for that sub-task and is followed by messages exchanged between the human and machine with message-tags by the interaction model described in section 2.1
Task Description: Title: Predicting star formation properties. Background: The ultimate goal is to determine a few output parameters such as the current star formation rate, the stellar mass, dust mass. The input dataset located at ’data/filtered_galaxies.csv’ consists of 22 input parameters and 3 output parameters. The data columns are: [’FUV_flux’, ’NUV_flux’, ’u_flux’, ’g_flux’, ’r_flux’, ’i_flux’, ’z_flux’, ’X_flux’, ’Y_flux’, ’J_flux’, ’H_flux’, ’K_flux’, ’W1_flux’, ’W2_flux’, ’W3_flux’, ’W4_flux’, ’P100_flux’, ’P160_flux’, ’S250_flux’, ’S350_flux’, ’S500_flux’, ’Z’, ’SFR_0_1Gyr_percentile50’, ’L_dust_percentile50’, ’mass_stellar_percentile50’]. The three output parameters that need to be predicted are: ’SFR_0_1Gyr_percentile50’, ’L_dust_percentile50’, ’mass_stellar_percentile50’. What needs to be done:Use the flux measurements (inputs) to predict the outputs (’SFR_0_1Gyr_percentile50’, ’L_dust_percentile50’, ’mass_stellar_percentile50’. For evaluating the results, use a custom error metric: Expected output: ML model(s) that can predict the star formation properties with minimal error.
Session 1: Perform exploratory data analysis Description: load data and do exploratory data analysis. Pre-condition: data/filtered_galaxies.csv Post-condition: graphs plotted and better understanding of the data. Data loaded as df.
Session 2: Feature engineering Description: Feature engineering. Normalise the data by removing the mean and scaling to unit variance. Split the data into training and test sets. Pre-condition: the data is loaded in the variable df Post-condition: Normalised train and test data to be given to ML model to train.
Session 3: Train ML models Description: Train a ML/AI model to predict the three output parameters. I leave it upto you based on your expertise in astrophysics and machine learning, whether to build a single model to predict all the three parameters or to build separate models for each. I need a model that will result in the least RMSE and the custom error given. Monitor the time taken to train the models for each output parameters. Save the model in a format that can be loaded again easily to compute various evaluation parameters. Save the predicted values and the actual values in a csv file for each of the three output parameters. Pre-condition: x_train, y_train, x_test, y_test datasets are available. Post-condition: Model trained and saved. Predicted and actual values saved.
Session 4: Reporting on evaluation metric and validation plots Description: Load the file with actual and predicted values saved. Evaluate the predictions with the actual values and report the metric. Report on the custom error, root mean square error, mean absolute deviation, r2 adjusted. Create plots of actual vs predicted values. Make the plot look sophisticated and professional. Pre-condition: actual and predicted values are stored as ’y_test,y_pred’ in: f’{output}_actual_vs_predicted.csv’ for outputs = [’SFR_0_1Gyr_percentile50’, ’L_dust_percentile50’, ’mass_stellar_percentile50’]. Post-condition: evaluation metric reported and saved on the custom error metric. Plots created and saved.